OpenStudy (anonymous):

what are the solutions to the following system of equations? y = x2 + 3x - 7 3x - y = -2

mathslover (mathslover):

Put the value of y in the second equation.

mathslover (mathslover):

From equation 1 : \( y = x^2 + 3x - 7 \) Now, put this value of y at the place of y in the second equation. \(3x - y = -2 \\ 3x - (x^2 + 3x - 7 ) = -2 \) Now, you can simplify this and obtain and equaton in single variable. Solve for x and then plug-in the value of x in the equation 1. Solve for y then!